How to solve -4/3x+4/9=-x^2
Answer by stanbon(75887)

solve -4/3x+4/9=-x^2
---
Multiply thru by -9 go get:
12x -4 = 9x^2
----
9x^2 - 12x + 4 = 0
----
Factor:
(3x-2)(3x-2) = 0
--------------------------
x = 2/3 with multiplicity 2.
==============
Cheers,
Stan H.
==============